OpenStudy (anonymous):

Fill in the blank. If necessary, use the slash mark ( / ) for a fraction bar. If sin0 = 3/5, then tan= _____.

OpenStudy (anonymous):

sin theta = 3/5 theta= 36.86 tan 36.86 = .75 = 3/4

OpenStudy (anonymous):

Thanls

OpenStudy (anonymous):

so do i put 37

OpenStudy (anonymous):

or 3/4

OpenStudy (anonymous):

3/4 :))

OpenStudy (anonymous):

okay<33